# Function spaces for decoupling

### v2

@@ 3.2. Wave packet transforms
endowed with the measure $\mathrm{d}x\mathrm{d}\omega\frac{\mathrm{d}\sigma}{\sigma}$. Note that the map $(x,\omega,\sigma)\mapsto(x,\sigma^{-1}\omega)$ identifies $S^{*}_{+}\mathbb{R}^{n}$ with $\mathbb{R}^{n}\times(\mathbb{R}^{n}\setminus\{0\})$, i.e. phase space minus the zero section.
To properly define our wave packet transforms, it will be convenient to work with a class of test functions on $S^{*}_{+}\mathbb{R}^{n}$ and the associated distributions. As in [19, Section 2.2], we let $\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ consist of those $G\in L^{\infty}(S^{*}_{+}\mathbb{R}^{n})$ such that
$(x,\omega,\sigma)\mapsto(1+|x|+\max(\sigma,\sigma^{-1}))^{N}G(x,\omega,\sigma)$
is an element of $L^{\infty}(S^{*}_{+}\mathbb{R}^{n})$ for all $N\geq 0$, endowed with the topology generated by the corresponding weighted $L^{\infty}$ norms. Let $\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})$ be the space of continuous linear $F:\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})\to{\mathbb{C}}$, endowed with the topology induced by $\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$. We denote the duality between $F\in\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})$ and $G\in\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ by $\langle F,G\rangle_{S^{*}_{+}\mathbb{R}^{n}}$. If $F\in L^{1}_{\mathrm{loc}}(S^{*}_{+}\mathbb{R}^{n})$ is such that
$G\mapsto\int_{S^{*}_{+}\mathbb{R}^{n}}F(x,\omega,\sigma)G(x,\omega,\sigma)\mathrm{d}x\mathrm{d}\omega\frac{\mathrm{d}\sigma}{\sigma}$
@@
on $S^{*}_{+}\mathbb{R}^{n}$, consisting of all $F\in L^{1}_{\mathrm{loc}}(S^{*}\mathbb{R}^{n})$ such that
@@
It is straightforward to check (see also [19, Lemma 2.10]) that
@@
(3.9) $VF(x):=$ $\int_{0}^{1}\int_{S^{n-1}}\psi_{\nu,\tau}(D)F(\cdot,\nu,\tau)(x)\mathrm{d}\nu\frac{\mathrm{d}\tau}{\tau}$
@@
    $W:L^{2}(\mathbb{R}^{n})\to L^{2}(S^{*}_{+}\mathbb{R}^{n})$ is an isometry;
@@
    $W:\mathcal{S}(\mathbb{R}^{n})\to\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ and $W:\mathcal{S}^{\prime}(\mathbb{R}^{n})\to\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})$ are continuous;
@@
    $V:\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})\to\mathcal{S}(\mathbb{R}^{n})$ is continuous, and $V$ extends uniquely to a continuous map $V:\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})\to\mathcal{S}^{\prime}(\mathbb{R}^{n})$;
@@
The statement and the proof are essentially contained in [19, Lemma 4.3]. More precisely, (1) follows from the definition of $\rho_{0}$ in (3.4), given that we have assumed that the surface measure $\mathrm{d}\omega$ on $S^{n-1}$ is unit normalized. Moreover, it is proved in [19, Lemma 4.3] that $W:\mathcal{S}(\mathbb{R}^{n})\to\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ and $V:\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})\to\mathcal{S}(\mathbb{R}^{n})$ are continuous. The remaining statements follow upon confirming that $\langle Wf,G\rangle_{S^{*}_{+}\mathbb{R}^{n}}=\langle f,VG\rangle_{\mathbb{R}^{n}}$ for all $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ and $G\in\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$. ∎
@@ 3.3. Operators on phase space
(3.12) $\widetilde{V}F(x):=$ $\int_{0}^{1}\int_{S^{n-1}}\widetilde{\psi}_{\nu,\tau}(D)F(\cdot,\nu,\tau)(x)\mathrm{d}\nu\frac{\mathrm{d}\tau}{\tau}$
@@
$W\widetilde{V}:L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})\to L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$
@@
$\lesssim\sigma^{\frac{n-1}{4}}\int_{1/4}^{4}\Big{(}\int_{\mathbb{R}^{n}}\Big{(}\int_{0}^{\infty}|F_{I}(x,\omega,\alpha\sigma)|^{2}\frac{\mathrm{d}\sigma}{\sigma}\Big{)}^{p/2}\mathrm{d}x\Big{)}^{1/p}\frac{\mathrm{d}\alpha}{\alpha}$
@@
Since $p\in(1,\infty)$, the space $X:=L^{p}(\mathbb{R}^{n};L^{2}(\mathbb{R}_{+}))$ is a Banach lattice with the UMD property (see e.g. [24, Section 4.2.c]). Given that we have $q\in(1,\infty)$ as well, [31, Theorem 3] then implies that $M_{S^{n-1}}$ is bounded on $L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})=L^{q}(S^{n-1};X)$. We now find
@@ 4.2. Connection to wave packet transforms
Let $p,q\in(1,\infty)$. Then $W:\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})\to L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$ is an isomorphism onto a complemented subspace, $V:L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})\to\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is bounded, and
(4.3) $V:L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})/\ker(V)\to\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$
@@
The first statement follows from Proposition 4.3, Theorem 3.4 and Proposition 3.3 (5), since $WV$ is a bounded projection on $L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$ with image $W\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$, and the kernel of this projection furnishes a complementary subspace. Next, note that $VF\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ for each $F\in L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$, by (3.7) and Proposition 3.3 (3). Hence the second statement also follows from Proposition 4.3 and Theorem 3.4. For the same reasons, $V:L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})\to\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is surjective, which implies (4.3). The final statement in turn follows from (4.3), or alternatively because $W:\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})\to W\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is an isomorphism onto a complemented subspace. ∎
@@ 4.3. Basic properties
Firstly, since $\langle D\rangle^{s}:\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})\to\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is an isomorphism, we may suppose throughout that $s=0$.
@@
We also claim that, if $(f_{k})_{k=0}^{\infty}$ is a uniformly bounded sequence in $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$, and if $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ is such that $f_{k}\to f$ in $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ as $k\to\infty$, then $f\in\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$. To prove this claim, recall from Lemma 3.2 that $\psi_{\omega,\sigma}\in C^{\infty}_{c}(\mathbb{R}^{n})$ for all $\omega\in S^{n-1}$ and $\sigma>0$. The same is true for the functions $\phi_{\omega,\sigma}$ defined by
@@
$\Phi(\sigma D)\varphi_{\omega}(D)f_{k}(x)=\mathcal{F}^{-1}(\phi_{\omega,\sigma})\ast f_{k}(x)\to\mathcal{F}^{-1}(\phi_{\omega,\sigma})\ast f(x)=\Phi(\sigma D)\varphi_{\omega}(D)f(x)$
@@
$|\Phi(\sigma D)\varphi_{\omega}(D)f(x)|=\lim_{k\to\infty}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}(x)|\leq\liminf_{k\to\infty}\sup_{\tau>0}|\Phi(\tau D)\varphi_{\omega}(D)f_{k}(x)|,$
@@
$\int_{\mathbb{R}^{n}}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f(x)|^{p}\mathrm{d}x$ $\leq\int_{\mathbb{R}^{n}}\liminf_{k\to\infty}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}(x)|^{p}\mathrm{d}x$
$\leq\liminf_{k\to\infty}\int_{\mathbb{R}^{n}}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}(x)|^{p}\mathrm{d}x.$
@@
$\leq\int_{S^{n-1}}\liminf_{k\to\infty}\big{\|}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}|\big{\|}_{L^{p}(\mathbb{R}^{n})}^{q}\mathrm{d}\omega$
$\leq\liminf_{k\to\infty}\int_{S^{n-1}}\big{\|}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}|\big{\|}_{L^{p}(\mathbb{R}^{n})}^{q}\mathrm{d}\omega\leq\liminf_{k\to\infty}\|f_{k}\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}^{q}<\infty,$
@@
After this preliminary work, let $(f_{k})_{k=0}^{\infty}$ be a Cauchy sequence in $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$. Then, by what we have just shown, there exists an $f\in\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ such that $f_{k}\to f$ in the topology of $\mathcal{S}^{\prime}(\mathbb{R}^{n})$. To see that in fact $\|f_{k}-f\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}\to 0$, one can argue as above. More precisely, for each $k\in\mathbb{Z}_{+}$, Fatou’s lemma yields
@@
$\leq\int_{S^{n-1}}\liminf_{l\to\infty}\big{\|}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)(f_{k}-f_{l})|\big{\|}_{L^{p}(\mathbb{R}^{n})}^{q}\mathrm{d}\omega$
$\leq\liminf_{l\to\infty}\int_{S^{n-1}}\big{\|}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)(f_{k}-f_{l})|\big{\|}_{L^{p}(\mathbb{R}^{n})}^{q}\mathrm{d}\omega$
$=\liminf_{l\to\infty}\|f_{k}-f_{l}\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}^{q}\leq\limsup_{l\to\infty}\|f_{k}-f_{l}\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}^{q}.$
@@
[…]
@@
[…]
@@
Finally, to see that $\mathcal{S}(\mathbb{R}^{n})$ in fact lies dense in $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$, let $f\in\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ be given. Then $Wf\in L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$, by Proposition 4.3. Since $\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})\subseteq L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$ lies dense, there exists a sequence $(F_{j})_{j=0}^{\infty}\subseteq\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ such that $F_{j}\to Wf$ in $L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$, as $j\to\infty$. Then, by Proposition 3.3 and Corollary 4.4, $(VF_{j})_{j=0}^{\infty}\subseteq\mathcal{S}(\mathbb{R}^{n})$ and
$\|VF_{j}-f\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}=\|VF_{j}-VWf\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}\lesssim\|F_{j}-Wf\|_{L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})}\to 0$
as $j\to\infty$. ∎

### v3

@@ 3.2. Wave packet transforms
endowed with the measure $\mathrm{d}x\mathrm{d}\omega\frac{\mathrm{d}\sigma}{\sigma}$. Note that the map $(x,\omega,\sigma)\mapstochar\rightarrow(x,\sigma^{-1}\omega)$ identifies $S^{*}_{+}\mathbb{R}^{n}$ with $\mathbb{R}^{n}\times(\mathbb{R}^{n}\setminus\{0\})$, i.e. phase space minus the zero section.
To properly define our wave packet transforms, it will be convenient to work with a class of test functions on $S^{*}_{+}\mathbb{R}^{n}$ and the associated distributions. As in [20, Section 2.2], we let $\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ consist of those $G\in L^{\infty}(S^{*}_{+}\mathbb{R}^{n})$ such that
$(x,\omega,\sigma)\mapstochar\rightarrow(1+|x|+\max(\sigma,\sigma^{-1}))^{N}G(x,\omega,\sigma)$
is an element of $L^{\infty}(S^{*}_{+}\mathbb{R}^{n})$ for all $N\geq 0$, endowed with the topology generated by the corresponding weighted $L^{\infty}$ norms. Let $\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})$ be the space of continuous linear $F:\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})\rightarrow{\mathbb{C}}$, endowed with the topology induced by $\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$. We denote the duality between $F\in\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})$ and $G\in\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ by $\langle F,G\rangle_{S^{*}_{+}\mathbb{R}^{n}}$. If $F\in L^{1}_{\mathrm{loc}}(S^{*}_{+}\mathbb{R}^{n})$ is such that
$G\mapstochar\rightarrow\int_{S^{*}_{+}\mathbb{R}^{n}}F(x,\omega,\sigma)G(x,\omega,\sigma)\mathrm{d}x\mathrm{d}\omega\frac{\mathrm{d}\sigma}{\sigma}$
@@
on $S^{*}_{+}\mathbb{R}^{n}$, consisting of all $F\in L^{1}_{\mathrm{loc}}(S^{*}_{+}\mathbb{R}^{n})$ such that
@@
It is straightforward to check (see also [20, Lemma 2.10]) that
@@
(3.9) $VF(x)$= $\int_{0}^{1}\int_{S^{n-1}}\psi_{\nu,\tau}(D)F(\cdot,\nu,\tau)(x)\mathrm{d}\nu\frac{\mathrm{d}\tau}{\tau}$
@@
    $W:L^{2}(\mathbb{R}^{n})\rightarrow L^{2}(S^{*}_{+}\mathbb{R}^{n})$ is an isometry;
@@
    $W:\mathcal{S}(\mathbb{R}^{n})\rightarrow\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ and $W:\mathcal{S}^{\prime}(\mathbb{R}^{n})\rightarrow\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})$ are continuous;
@@
    $V:\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})\rightarrow\mathcal{S}(\mathbb{R}^{n})$ is continuous, and $V$ extends uniquely to a continuous map $V:\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})\rightarrow\mathcal{S}^{\prime}(\mathbb{R}^{n})$;
@@
The statement and the proof are essentially contained in [20, Lemma 4.3]. More precisely, (1) follows from the definition of $\rho_{0}$ in (3.4), given that we have assumed that the surface measure $\mathrm{d}\omega$ on $S^{n-1}$ is unit normalized. Moreover, it is proved in [20, Lemma 4.3] that $W:\mathcal{S}(\mathbb{R}^{n})\rightarrow\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ and $V:\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})\rightarrow\mathcal{S}(\mathbb{R}^{n})$ are continuous. The remaining statements follow upon confirming that $\langle Wf,G\rangle_{S^{*}_{+}\mathbb{R}^{n}}=\langle f,VG\rangle_{\mathbb{R}^{n}}$ for all $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ and $G\in\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$. ∎
@@ 3.3. Operators on phase space
(3.12) $\widetilde{V}F(x)$= $\int_{0}^{1}\int_{S^{n-1}}\widetilde{\psi}_{\nu,\tau}(D)F(\cdot,\nu,\tau)(x)\mathrm{d}\nu\frac{\mathrm{d}\tau}{\tau}$
@@
$W\widetilde{V}:L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})\rightarrow L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$
@@
$\lesssim\int_{1/4}^{4}\Big(\int_{\mathbb{R}^{n}}\Big(\int_{0}^{\infty}\sigma^{\frac{n-1}{2}}|F_{I}(x,\omega,\alpha\sigma)|^{2}\frac{\mathrm{d}\sigma}{\sigma}\Big)^{p/2}\mathrm{d}x\Big)^{1/p}\frac{\mathrm{d}\alpha}{\alpha}$
@@
Since $p\in(1,\infty)$, the space $X:=L^{p}(\mathbb{R}^{n};L^{2}(\mathbb{R}_{+}))$ is a Banach lattice with the UMD property (see e.g. [26, Section 4.2.c]). Given that we have $q\in(1,\infty)$ as well, [40, Theorem 1.5] (see also [33]) then implies that $M_{S^{n-1}}$ is bounded on $L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})=L^{q}(S^{n-1};X)$. We now find
@@ 4.2. Connection to wave packet transforms
Let $p,q\in(1,\infty)$. Then $W:\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})\rightarrow L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$ is an isomorphism onto a complemented subspace, $V:L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})\rightarrow\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is bounded, and
(4.3) $V:L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})/\ker(V)\rightarrow\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$
@@
The first statement follows from Proposition 4.3, Theorem 3.4 and Proposition 3.3 (5), since $WV$ is a bounded projection on $L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$ with image $W\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$, and the kernel of this projection furnishes a complementary subspace. Next, note that $VF\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ for each $F\in L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$, by (3.7) and Proposition 3.3 (3). Hence the second statement also follows from Proposition 4.3 and Theorem 3.4. For the same reasons, $V:L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})\rightarrow\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is surjective, which implies (4.3). The final statement in turn follows from (4.3), or alternatively because $W:\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})\rightarrow W\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is an isomorphism onto a complemented subspace. ∎
@@ 4.3. Basic properties
Firstly, since $\langle D\rangle^{s}:\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})\rightarrow\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is an isomorphism, we may suppose throughout that $s=0$.
@@
We also claim that, if $(f_{k})_{k=0}^{\infty}$ is a uniformly bounded sequence in $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$, and if $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ is such that $f_{k}\rightarrow f$ in $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ as $k\rightarrow\infty$, then $f\in\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$. To prove this claim, recall from Lemma 3.2 that $\psi_{\omega,\sigma}\in C^{\infty}_{c}(\mathbb{R}^{n})$ for all $\omega\in S^{n-1}$ and $\sigma>0$. The same is true for the functions $\phi_{\omega,\sigma}$ defined by
@@
$\Phi(\sigma D)\varphi_{\omega}(D)f_{k}(x)=\mathcal{F}^{-1}(\phi_{\omega,\sigma})\ast f_{k}(x)\rightarrow\mathcal{F}^{-1}(\phi_{\omega,\sigma})\ast f(x)=\Phi(\sigma D)\varphi_{\omega}(D)f(x)$
@@
$|\Phi(\sigma D)\varphi_{\omega}(D)f(x)|=\lim_{k\rightarrow\infty}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}(x)|\leq\liminf_{k\rightarrow\infty}\sup_{\tau>0}|\Phi(\tau D)\varphi_{\omega}(D)f_{k}(x)|,$
@@
$\int_{\mathbb{R}^{n}}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f(x)|^{p}\mathrm{d}x$ $\leq\int_{\mathbb{R}^{n}}\liminf_{k\rightarrow\infty}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}(x)|^{p}\mathrm{d}x$
$\leq\liminf_{k\rightarrow\infty}\int_{\mathbb{R}^{n}}\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}(x)|^{p}\mathrm{d}x.$
@@
$\leq\int_{S^{n-1}}\liminf_{k\rightarrow\infty}\big\|\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}|\big\|_{L^{p}(\mathbb{R}^{n})}^{q}\mathrm{d}\omega$
$\leq\liminf_{k\rightarrow\infty}\int_{S^{n-1}}\big\|\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)f_{k}|\big\|_{L^{p}(\mathbb{R}^{n})}^{q}\mathrm{d}\omega\leq\liminf_{k\rightarrow\infty}\|f_{k}\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}^{q}<\infty,$
@@
After this preliminary work, let $(f_{k})_{k=0}^{\infty}$ be a Cauchy sequence in $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$. Then, by what we have just shown, there exists an $f\in\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ such that $f_{k}\rightarrow f$ in the topology of $\mathcal{S}^{\prime}(\mathbb{R}^{n})$. To see that in fact $\|f_{k}-f\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}\rightarrow 0$, one can argue as above. More precisely, for each $k\in\mathbb{Z}_{+}$, Fatou’s lemma yields
@@
$\leq\int_{S^{n-1}}\liminf_{l\rightarrow\infty}\big\|\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)(f_{k}-f_{l})|\big\|_{L^{p}(\mathbb{R}^{n})}^{q}\mathrm{d}\omega$
$\leq\liminf_{l\rightarrow\infty}\int_{S^{n-1}}\big\|\sup_{\sigma>0}|\Phi(\sigma D)\varphi_{\omega}(D)(f_{k}-f_{l})|\big\|_{L^{p}(\mathbb{R}^{n})}^{q}\mathrm{d}\omega$
$=\liminf_{l\rightarrow\infty}\|f_{k}-f_{l}\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}^{q}\leq\limsup_{l\rightarrow\infty}\|f_{k}-f_{l}\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}^{q}.$
@@
$\limsup_{k\rightarrow\infty}\|f_{k}-f\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}\leq\limsup_{k\rightarrow\infty}\limsup_{l\rightarrow\infty}\|f_{k}-f_{l}\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}=0.\qed$
@@
$\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})\xrightarrow{W}L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})\subseteq\mathcal{J}^{\prime}(S^{*}_{+}\mathbb{R}^{n})\xrightarrow{V}\mathcal{S}^{\prime}(\mathbb{R}^{n}),$
@@
Finally, to see that $\mathcal{S}(\mathbb{R}^{n})$ in fact lies dense in $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$, let $f\in\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ be given. Then $Wf\in L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$, by Proposition 4.3. Since $\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})\subseteq L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$ lies dense, there exists a sequence $(F_{j})_{j=0}^{\infty}\subseteq\mathcal{J}(S^{*}_{+}\mathbb{R}^{n})$ such that $F_{j}\rightarrow Wf$ in $L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$, as $j\rightarrow\infty$. Then, by Proposition 3.3 and Corollary 4.4, $(VF_{j})_{j=0}^{\infty}\subseteq\mathcal{S}(\mathbb{R}^{n})$ and
$\|VF_{j}-f\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}=\|VF_{j}-VWf\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})}\lesssim\|F_{j}-Wf\|_{L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})}\rightarrow 0$
as $j\rightarrow\infty$. ∎